# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v3

@@ 14. Feature C bucket 5 - Binary
Figure 6. Prompt passed to the LLM to generate reward function based on the problem scenarios in the Synthetic Domain.
@@ 11. Education level 7/7 – Post graduate - Binary
Figure 7. Prompt passed to the LLM to generate reward function based on the problem scenarios in the Real World Domain.
@@
Figure 8. Prompt passed to the LLM to choose a reward function based on the context of problem scenario in the Synthetic Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 9. Prompt passed to the LLM to choose a reward function based on the context of problem scenario in Real World Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 10. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
Figure 11. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
Figure 12. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
Figure 13. Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
Figure 14. Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.
@@
Figure 15. Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.

### v4

@@ 14. Feature C bucket 5 - Binary
Figure 7: Prompt passed to the LLM to generate reward function based on the problem scenarios in the Synthetic Domain.
@@ 11. Education level 7/7 – Post graduate - Binary
Figure 8: Prompt passed to the LLM to generate reward function based on the problem scenarios in the Real World Domain.
@@
Figure 9: Prompt passed to the LLM to choose a reward function based on the context of problem scenario in the Synthetic Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 10: Prompt passed to the LLM to choose a reward function based on the context of problem scenario in Real World Domain, the generated reward functions and the reward distribution corresponding to every reward function.
@@
Figure 11: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
Figure 12: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional examples on what to look at when choosing a reward function aligned with the preference.
@@
Figure 13: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
Figure 14: Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to minimize the unintended utility shifts in dimensions not specified in the preference.
@@
Figure 15: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.
@@
Figure 16: Continued: Enhanced Prompt passed to the LLM to choose a reward function based on the context of the problem scenario in the Real World Domain, the generated reward functions, the reward distributions corresponding to every reward function and additional information to maximize the overall utility.